A7. A recent study shows that the annual cost of maintaining a building in Accra averages 200 Cedis with a variance of 260 Cedis. If a tax of 30% is introduced on all items associated with the maintenance of building i.e. everything is made 30% more expensive. Calculate the standard deviation of the annual cost of maintaining a building in Accra
1
Expert's answer
2020-05-12T18:43:14-0400
"V(X)=\sigma^2=\sum(x_i-\mu)^2p(x_i)"

"V(aX)=\sigma_{aX}^2=a^2 V(X)"

"V(1.3X)=1.3^2V(X)=1.69(260)=439.4"

"\sigma_{1.3X}=\sqrt{V(1.3X)}=\sqrt{439.4}\approx20.96"

The standard deviation of the annual cost of maintaining a building in Accra is "20.96."